OpenStudy (anonymous):

Ralph is 3 times as old as Sara. In 4 years, Ralph will be only twice as old as Sara will be then. Find Ralph's age now.

OpenStudy (kainui):

Try to write out some of this as an equation. Maybe simplify Ralph's age as making it a variable like R and Sara's age a variable like S.

OpenStudy (kainui):

You can think that when it says Ralph is 3 times as old as Sara that means we can replace Ralph with R, "is" with an equals sign, times with a multiplication sign, and Sara with an S and you have one equation made already no sweat.

OpenStudy (anonymous):

It confuses me :(

OpenStudy (shamim):

So R=3s

OpenStudy (kainui):

Well I'm here to help. Try stuff out, don't worry about making mistakes because if someone gives you the answer, that person can't give you the answer on the test and you'll be left high and dry.

OpenStudy (anonymous):

ya then after 4 years r+4=2s 2s-r=4

OpenStudy (anonymous):

now solve it with3s-r=0

OpenStudy (kainui):

@rathanreddy you're kind of wasting my time when you do that.

OpenStudy (anonymous):

k then utilize your time

OpenStudy (blurbendy):

lol

OpenStudy (nurali):

let Sara's age = x. Then Ralph's age now = 3x. 3x+4= 2(x+4) 3x+4 =2x+8 x =4 (Sara's age now) Ralph is 3x=12 now